# Diederich–Fornæss index and global regularity in the $\overline{\partial}$–Neumann problem: domains with comparable Levi eigenvalues

### v4

@@ 1. Introduction
The question what the implications of the Diederich–Fornæss index are for global regularity arose in earnest after the results on the worm domains in [7, 2, 14] and after [9], where the authors showed that if $\Omega$ admits a defining function that is plurisubharmonic at the boundary, then the Bergman projections and $\overline{\partial}$–Neumann operators (at all form levels) on $\Omega$ are globally exactly regular, i.e. are continuous in Sobolev–$s$ norms for $s\geq 0$. A quantitative study of this question was initiated in [29] where it was shown that if the index is one, and there is some control on the defining functions $\rho_{\eta}$ as $\eta\rightarrow 1^{-}$, then global regularity holds. More precisely, the condition is that $\liminf_{(\eta\rightarrow 1^{-})}((1-\eta)^{1/3}\max_{b\Omega}|\nabla h_{\eta}|)=0$, where $h_{\eta}$ is given by $\rho_{\eta}=e^{h_{\eta}}\rho$ for some fixed defining function $\rho$. The exponent $1/3$ was improved to $1/2$ in [23]. In [6], it is shown that if the index is $\eta_{0}\leq 1$, then regularity holds for $0\leq s<\eta_{0}/2\;(\leq 1/2)$. In [33], the authors, among other things, generalize the ideas from [29] to $q$–convex domains, with a corresponding notion of the index. More recently, the first author showed in [32] that index one implies global regularity for domains with comparable eigenvalues of the Levi form without assuming any control on the defining functions $\rho_{\eta}$, but instead making a technical assumption that controls the geometry of the set of infinite type boundary points.
@@ 2. Notation and Preliminaries
We use the following notation for weighted Sobolev norms: for a real valued function $g$ so that $e^{-g}\in C^{\infty}(\overline{\Omega})$, we set $\|u\|_{k,g}=\left(\sum_{s=0}^{k}\int_{\Omega}|\nabla^{s}u|^{2}e^{-g}\right)^{1/2}$, where $\nabla^{s}$ denotes the vector of all derivatives of order $s$. If $u$ is a form, the derivatives act coefficients as usual (in Euclidean coordinates, unless stated otherwise).
@@
([36], (5.84)).
The forms $\alpha_{h}$ play a crucial role for Sobolev estimates in the $\overline{\partial}$–Neumann problem in that they control commutators of certain vector fields with $\overline{\partial}$ and $\overline{\partial}^{*}$(compare (4) above, [36], [25],[15]). We recall two lemmas from [25] that give explicit formulas for these commutators, modulo benign errors. These formulas are stated for forms supported in a special boundary chart, with the usual orthonormal boundary frame $L_{1},\cdots,L_{n}$ and dual frame $\omega_{1},\cdots,\omega_{n}$. Vector fields (differential operators) on forms are acting coefficient wise in this special chart. Let $v=\sideset{}{{}^{\prime}}{\sum}_{J}v_{J}\overline{\omega_{J}}\in C^{\infty}_{(0,q)}(\overline{\Omega})$, and let $0\leq q\leq n$. In our notation, and with our normalizations, the formulas from [25] are as follows.
@@
(7) [¯$[\overline{\partial}^{*},(e^{-h}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}]v\\ =-k\sum_{j<n}\sideset{}{{}^{\prime}}{\sum}_{S}(L_{j}h+\alpha^{\rho}(L_{j}))((e^{-h}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}v_{jS})\;\overline{\omega_{S}}+A_{h}(v)+B_{h}(v)+C_{h}(v_{N}),$
where $A_{h}(v)$ and $B_{h}(v)$ are of the same type as in Lemma 3 (but different operators), and $C_{h}(v_{N})$ is of order $k$ on the normal component $v_{N}=\sideset{}{{}^{\prime}}{\sum}_{S}v_{Sn}\;(\overline{\omega_{S}}\wedge\overline{\omega_{n}})$ of $v$.
Note that when $j<n$, then $\overline{L_{j}}h+\alpha^{\rho}(\overline{L_{j}})$ equals $\alpha_{h}(\overline{L_{j}})$ (in view of (5)). Likewise, $L_{j}h+\alpha^{\rho}(L_{j})=\alpha_{h}(L_{j})$ when $j<n$. Thus the form $\alpha_{h}$ controls the commutators in (6) and (7), except for the term with $j=n$ in (6). However, it has been understood at least since [9] that control of this term comes for free, see (39) below. The error terms $A_{h}(v)$, $B_{h}(v)$, and $C_{h}(v_{N})$ are normally under control ($B_{h}(v)$ is lower order, for $A_{h}(v)$ use Lemma 1, and the normal component acts as a subelliptic multiplier, see estimate (2.96) in [36]; Proposition 4.7, part (G), in [28]). In fact, in [25] the error terms are given only in this estimated form. On the other hand, for the proof of the lemmas, which results from computing the case $k=1$ in the boundary frame, and then using induction on $k$, it is convenient to have an expression for the terms. The factor $k$ on the right hand sides of (6) and (7) arises in the induction from the commutator formula $[A,T^{k}]=\sum_{j=1}^{k}\binom{k}{j}[\cdots[[A,T],T]\cdots]\,T^{k-j}=k\,[[A,T],T]T^{k-1}$ $+$ terms of order at most $(k-1)$ (see [18], Lemma 2 on page 418, or [36], formula (3.54)).
@@ 3. Diederich-Fornæss index and estimates on D’Angelo forms
(8) $DF(\Omega)=\sup\left\{0<\eta<1:\left(\frac{\eta}{1-\eta}(\omega_{\rho}\wedge\overline{\omega_{\rho}})-\overline{\partial}\omega_{\rho}\right)(L\wedge\overline{L})\leq 0\;;\;p\in\Sigma\;,\;L\in\mathcal{N}_{p}\right\}\;,$
where $\Sigma\subset b\Omega$ denotes the set of weakly pseudoconvex boundary points, $\mathcal{N}_{p}$ denotes the null space of the Levi form at $p\in b\Omega$, and the supremum is over all $\eta$ so that there exists a defining function $\rho_{\eta}$ such that the inequality holds. Because $|\alpha_{\eta}(L)|^{2}=(\omega_{\eta}\wedge\overline{\omega_{\eta}})(L\wedge\overline{L})$ and ¯$\overline{\partial}\omega_{\eta}(L\wedge\overline{L})=\overline{\partial}\alpha_{\eta}(L\wedge\overline{L})$, (8) implies $|\alpha_{\eta}(L(p))|^{2}\lesssim(1-\eta)\overline{\partial}\alpha_{\eta}(L(p)\wedge\overline{L(p)})\;;\;p\in\Sigma\;,\;L(p)\in\mathcal{N}_{p}$. Computing $\alpha_{\eta}$ from its definition as a Lie derivative (see section 5.9 in [36]) shows that $\alpha_{\eta}=\alpha^{\rho}+dh_{\eta}-(Th_{\eta})\sigma^{\rho}$. Therefore ¯$\overline{\partial}\alpha_{\eta}=\overline{\partial}\alpha^{\rho}+\overline{\partial}dh_{\eta}-\overline{\partial}((Th_{\eta})\sigma^{\rho})$. The last term vanishes on $\mathcal{N}_{p}$: ¯$\overline{\partial}((Th_{\eta})\sigma^{\rho})(L\wedge\overline{L})=(\overline{\partial}(Th_{\eta})\wedge\sigma^{\rho})(L\wedge\overline{L})+(1/2)(Th_{\eta})\overline{\partial}(\partial\rho-\overline{\partial}\rho)(L\wedge\overline{L})$; the first term vanishes because $\sigma$ annihilates $L$ and $\overline{L}$, the second because ¯$\overline{\partial}\partial\rho(L\wedge\overline{L})=0$ when $L\in\mathcal{N}_{p}$. Since also ¯$\overline{\partial}dh_{\eta}=\overline{\partial}\partial h_{\eta}=-\partial\overline{\partial}h_{\eta}$, we have, for some constant $C$ that does not depend on $\eta$,
@@
Assumptions as in Theorem 1, $q_{0}\leq q\leq(n-1)$. There are a constant $C$ and, for $(1-\eta)$ small enough, a constant $C_{\eta}$ and a relatively compact subdomain $\Omega_{\eta}$ such that
@@ 4. Proof of Theorem 1
As usual, the arguments will involve absorbing terms, and so one has to know that these terms are finite. In order to insure that, we work with the regularized operators $N_{\delta,q}$ that result from elliptic regularization. That is, $N_{\delta,q}$ is the inverse of the Salvadorian operator $\Box_{\delta,q}$ associated with the quadratic form $Q_{\delta,q}(u,u)=\|\overline{\partial}u\|^{2}+\|\overline{\partial}^{*}u\|^{2}+\delta\|\nabla u\|^{2}$, with form domain $W^{1}_{(0,q)}(\Omega)\cap dom(\overline{\partial}^{*})$, where $\nabla$ denotes the vector of all first order derivatives of all coefficients of $u$. $N_{\delta,q}$ maps $L^{2}_{(0,q)}(\Omega)$ continuously into the form domain, endowed with the norm $Q_{\delta,q}(u,u)^{1/2}$. We will prove estimates with constants that are uniform in $\delta$; letting $\delta\rightarrow 0$ then gives the desired estimates for $N_{q}$ ([13], pages 102–103 and [36], section 3.3).
So $q$ is now fixed, $q_{0}\leq q$, and $N_{q+1}$, hence $P_{q}$, satisfy Sobolev estimates. We want to prove Sobolev estimates for $N_{\delta,q}$, with constants uniform in $\delta$ ($\delta$ small). Because $P_{q}$ is regular in Sobolev norms, Lemma 7 in [35] says that for all $s>0$, there is a constant $C_{s}$ independent of $\delta$ such that we have the estimate
@@
For $\eta\in(0,1)$, set $X_{\eta}=e^{-h_{\eta}+\rho g_{\eta}}L_{n}^{\rho}$, where $g_{\eta}$ is a smooth function that agrees on the boundary with $\alpha_{\eta}(\overline{L_{n}})/\overline{L_{n}}\rho$. The reason for this choice of exponent (rather than just $-h_{\eta}$) is that the terms where $j=n$ in the commutator (40) below will then vanish on the boundary. This fact is needed in the argument. Being able to make this choice is a manifestation of the principle that ‘commutator conditions in the normal (and any strictly pseudoconvex) direction come for free’ ([36], Section 5.7, [9], proof of the lemma). Note that because $g_{\eta}$ is only prescribed on the boundary, we may assume that $|\rho g_{\eta}|\leq 1/2k$. As usual, we need tangential differential operators to preserve the domain of $\overline{\partial}^{*}$, and so we let them act in special boundary charts ([36], section 2.2). The error this introduces is one order lower than the operator, and as a result, our estimates are not affected.
@@
The $\|u\|_{k}^{2}$ term is needed because of the compactly supported term from the partition of unity used for the special boundary charts; this term is not covered by Lemma 3. The estimate follows from interior elliptic regularity of $\Box_{\delta,q}$, with constants that are uniform in $\delta$. This uniformity holds because for $u\in dom(\Box_{\delta,q})$, $\Box_{\delta,q}u=-(1/4+\delta)\Delta u$, where $\Delta$ acts coefficient wise (see [36], formula (3.22)). We have used that when there is a barred derivative, or a complex tangential derivative, we can always commute it so that it acts first. Then there is no issue with the requirement that a form must be in the domain of $\overline{\partial}^{*}$ in order for Lemma 1 to apply. The error term this makes is of order at most $(k-1)$. Modulo terms controlled by ¯$\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}$, we have $\|(e^{-h_{\eta}}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}\lesssim\|e^{-kh_{\eta}}(L_{n}^{\rho}-\overline{L_{n}^{\rho}})^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}\leq e\|(X_{\eta}-\overline{X_{\eta}})^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}$ (since $2k|\rho g_{\eta}|\leq 1$). Upon also estimating the last term in (30) via a s.c.–l.c argument and absorbing the term $s.c\|\overline{\partial}^{*}N_{\delta,q}u\|_{k,2h_{\eta}}^{2}$ into the left hand side of (30), we thus arrive at
@@
Here, $\vartheta$ is the formal adjoint of $\overline{\partial}$, $(\partial/\partial\nu)$ denotes the normal derivative acting coefficient wise on forms, and $\omega_{n}$ is the form dual to the unit normal $L_{n}$.
@@
More precisely, it is shown in [35] that for $\eta$ given, there is $\delta_{0}(\eta)$ such that all the error terms in (31)–(33) are acceptable for (29) (for $l=k$), or can be absorbed into the left hand side of (29) when $\delta\leq\delta_{0}(\eta)$. We will use acceptable somewhat loosely for terms in an estimate that are either bounded appropriately by the right hand side, or can be absorbed into the left hand side. Note that upon adding (34) through (36), using Cauchy–Schwarz on each inner product, followed by a s.c.–l.c. argument, the squares of the norms of the left hand sides in (34)–(36) can be absorbed into the first term on the right hand side of the sum of (31)–(33). Therefore, in order to estimate the sum of the left hand sides of (31)–(33), we only have to estimate
@@
For the commutators with $\overline{\partial}$ and $\overline{\partial}^{*}$ we use Lemmas 3 and 4, with $h=h_{\eta}-\rho g_{\eta}$ and $N_{\delta,q}u$. It suffices to estimate the terms in (38) with $N_{\delta,q}u$ replaced by $\chi N_{\delta,q}u$, where $\chi\in C^{\infty}(\mathbb{C}^{n})$ is supported in a boundary chart. Via a suitable partition of unity (independent of $\eta$), these estimates will then be summed to obtain global estimates. Observe that near the boundary
(39) $\overline{L_{j}}h+\alpha^{\rho}(\overline{L_{j}})=-\overline{L_{j}}(\rho g_{\eta})+\overline{L_{j}}h_{\eta}+\alpha^{\rho}(\overline{L_{j}})=-\overline{L_{j}}(\rho g_{\eta})+\alpha_{\eta}(\overline{L_{j}})\;;\;j<n\;;$
the last equality follows from (5). Inserting this equality into Lemma 3 and estimating the error terms with Lemma 1 gives
@@
(40) [¯$\|[\overline{\partial},(X_{\eta}-\overline{X_{\eta}})^{k}]\chi N_{\delta,q}u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j\notin J}\int_{\Omega}|\overline{L_{j}}(\rho g_{\eta})-\alpha_{\eta}(\overline{L_{j}})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\;\;\;\;\;\;\;\;\;\;\;\;\\ +C_{\eta}\,(\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}+\|u\|_{k-2}^{2})\;.$
We only have to sum over $j\notin J$ because for $j\in J$, the term $\overline{\omega_{j}}\wedge\overline{\omega_{J}}$ vanishes. In the error terms, we omit the factor $\chi$ and use that the commutators of $\chi$ with $\overline{\partial}$ and $\overline{\partial}^{*}$ are of order zero. Because there are only finitely many $\chi$s, we may keep $C_{\eta}$ independent of $\chi$.
@@
The term $\|u\|_{k-2}^{2}$ again results from the compactly supported term that arises from letting $(X_{\eta}-\overline{X_{\eta}})$ act in special boundary charts; as above, the estimate follows from interior elliptic regularity of $\Box_{\delta,q}$, with constants that are uniform in $\delta$. The order $(k-1)$ error terms are acceptable by the induction hypothesis, and using (28) for $\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}$ shows that this term can be split into an absorbable term and a lower order term. From Lemma 4, we similarly obtain
@@
(41) [¯$\|[\overline{\partial}^{*},(X_{\eta}-\overline{X_{\eta}})^{k}]\chi N_{\delta,q}u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j\in J}\int_{\Omega}|L_{j}(\rho g_{\eta}-\alpha_{\eta}(L_{j})|^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;+C_{\eta}\,\left(\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}+\|(N_{\delta,q}u)_{N}\|_{k}^{2}+\|u\|_{k-2}^{2}\right)\;.$

### v5

@@ 1. Introduction
The question what the implications of the Diederich–Fornæss index are for global regularity arose in earnest after the results on the worm domains in [7, 2, 14] and after [9], where the authors showed that if $\Omega$ admits a defining function that is plurisubharmonic at the boundary, then the Bergman projections and $\overline{\partial}$–Neumann operators (at all form levels) on $\Omega$ are globally exactly regular, i.e., are continuous in Sobolev–$s$ norms for $s\geq 0$. A quantitative study of this question was initiated in [29] where it was shown that if the index is one, and there is some control on the defining functions $\rho_{\eta}$ as $\eta\rightarrow 1^{-}$, then global regularity holds. More precisely, the condition is that $\liminf_{(\eta\rightarrow 1^{-})}((1-\eta)^{1/3}\max_{b\Omega}|\nabla h_{\eta}|)=0$, where $h_{\eta}$ is given by $\rho_{\eta}=e^{h_{\eta}}\rho$ for some fixed defining function $\rho$. The exponent $1/3$ was improved to $1/2$ in [23]. In [6], it is shown that if the index is $\eta_{0}\leq 1$, then regularity holds for $0\leq s<\eta_{0}/2\;(\leq 1/2)$. In [33], the authors, among other things, generalize the ideas from [29] to $q$–convex domains, with a corresponding notion of the index. More recently, the first author showed in [32] that index one implies global regularity for domains with comparable eigenvalues of the Levi form without assuming any control on the defining functions $\rho_{\eta}$, but instead making a technical assumption that controls the geometry of the set of infinite type boundary points.
@@ 2. Notation and Preliminaries
We use the following notation for weighted Sobolev norms: for a real valued function $g$ so that $e^{-g}\in C^{\infty}(\overline{\Omega})$, we set $\|u\|_{k,g}=\left(\sum_{s=0}^{k}\int_{\Omega}|\nabla^{s}u|^{2}e^{-g}\right)^{1/2}$, where $\nabla^{s}$ denotes the vector of all derivatives of order $s$. If $u$ is a form, the derivatives act on coefficients as usual (in Euclidean coordinates, unless stated otherwise).
@@
([36], (5.84)). Also note that $\alpha^{\rho}(T^{\rho})=0$ (from the definition and the normalizations).
The forms $\alpha^{\rho}$ play a crucial role for Sobolev estimates in the $\overline{\partial}$–Neumann problem in that they control commutators of certain vector fields with $\overline{\partial}$ and $\overline{\partial}^{*}$(compare (4) above, [36], [25],[15]). We recall two lemmas from [25] that give explicit formulas for these commutators, modulo benign errors. These formulas are stated for forms supported in a special boundary chart, with the usual orthonormal boundary frame $L_{1},\cdots,L_{n}$ and dual frame $\omega_{1},\cdots,\omega_{n}$. Vector fields (differential operators) on forms are acting coefficient wise in this special chart. Let $v=\sideset{}{{}^{\prime}}{\sum}_{J}v_{J}\overline{\omega_{J}}\in C^{\infty}_{(0,q)}(\overline{\Omega})$, and let $0\leq q\leq n$. In our notation, and with our normalizations, the formulas from [25] are as follows.
@@
(7) [¯$[\overline{\partial}^{*},(e^{-h}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}]v\\ =-k\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{S}(L_{j}h+\alpha^{\rho}(L_{j}))((e^{-h}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}v_{jS})\;\overline{\omega_{S}}+A_{h}(v)+B_{h}(v),$
where $A_{h}(v)$ and $B_{h}(v)$ are of the same type as in Lemma 3 (but different operators).
Note that when $j<n$, then $\overline{L_{j}}h+\alpha^{\rho}(\overline{L_{j}})$ equals $\alpha_{h}(\overline{L_{j}})$ (in view of (5)). Likewise, $L_{j}h+\alpha^{\rho}(L_{j})=\alpha_{h}(L_{j})$ when $j<n$. Thus the form $\alpha_{h}$ controls the commutators in (6) and (7), except for the term with $j=n$ in (6). However, it has been understood at least since [9] that control of this term comes for free. The error terms $A_{h}(v)$, $B_{h}(v)$, and $C_{h}(v_{N})$ are normally under control ($B_{h}(v)$ is lower order, for $A_{h}(v)$ use Lemma 1, and the normal component acts as a subelliptic multiplier, see estimate (2.96) in [36]; Proposition 4.7, part (G), in [28]). In fact, in [25] the error terms are given only in this estimated form. On the other hand, for the proof of the lemmas, which results from computing the case $k=1$ in the boundary frame, and then using induction on $k$, it is convenient to have an expression for the terms. The factor $k$ on the right hand sides of (6) and (7) arises in the induction from the commutator formula $[A,T^{k}]=\sum_{j=1}^{k}\binom{k}{j}[\cdots[[A,T],T]\cdots]\,T^{k-j}=k\,[[A,T],T]T^{k-1}$ $+$ terms of order at most $(k-1)$ (see [18], Lemma 2 on page 418, or [36], formula (3.54)).
@@ 3. Diederich-Fornæss index and estimates on D’Angelo forms
(8) $DF(\Omega)=\sup\left\{0<\eta<1:\left(\frac{\eta}{1-\eta}(\omega_{\eta}\wedge\overline{\omega_{\eta}})-\overline{\partial}\omega_{\eta}\right)(L\wedge\overline{L})\leq 0\;;\;p\in\Sigma\;,\;L\in\mathcal{N}_{p}\right\}\;,$
where $\Sigma\subset b\Omega$ denotes the set of weakly pseudoconvex boundary points, $\mathcal{N}_{p}$ denotes the null space of the Levi form at $p\in b\Omega$, and the supremum is over all $\eta$ so that there exists a defining function $\rho_{\eta}$ such that the inequality holds. Because $|\alpha_{\eta}(L)|^{2}=(\omega_{\eta}\wedge\overline{\omega_{\eta}})(L\wedge\overline{L})$ and ¯$\overline{\partial}\omega_{\eta}(L\wedge\overline{L})=\overline{\partial}\alpha_{\eta}(L\wedge\overline{L})$, (8) implies $|\alpha_{\eta}(L(p))|^{2}\lesssim(1-\eta)\overline{\partial}\alpha_{\eta}(L(p)\wedge\overline{L(p)})\;;\;p\in\Sigma\;,\;L(p)\in\mathcal{N}_{p}$. Section 4.3 in [15] (see in particular Lemmas 4.6 and 4.7) gives that for $p\in\Sigma$, $L(p)\in\mathcal{N}_{p}$, we have ¯$\overline{\partial}\alpha_{\eta}(L(p)\wedge\overline{L(p)})=\overline{\partial}\alpha^{\rho}(L(p)\wedge\overline{L(p)})-\partial\overline{\partial}h_{\eta}(L(p)\wedge\overline{L_{p}})$. Consequently, we have, for some constant $C$ that does not depend on $\eta$,
@@
Assumptions as in Theorem 1, $q_{0}\leq q\leq(n-1)$. There are a constant $C$ and, for $(1-\eta)$ small enough, a constant $C_{\eta}$, such that
@@ 4. Proof of Theorem 1
As usual, the arguments will involve absorbing terms, and so one has to know that these terms are finite. In order to insure that, we work with the regularized operators $N_{\delta,q}$ that result from elliptic regularization. That is, $N_{\delta,q}$ is the inverse of the operator $\Box_{\delta,q}$ associated with the quadratic form $Q_{\delta,q}(u,u)=\|\overline{\partial}u\|^{2}+\|\overline{\partial}^{*}u\|^{2}+\delta\|\nabla u\|^{2}$, with form domain $W^{1}_{(0,q)}(\Omega)\cap dom(\overline{\partial}^{*})$, where $\nabla$ denotes the vector of all first order derivatives of all coefficients of $u$. $N_{\delta,q}$ maps $L^{2}_{(0,q)}(\Omega)$ continuously into the form domain, endowed with the norm $Q_{\delta,q}(u,u)^{1/2}$. We will prove estimates with constants that are uniform in $\delta$; letting $\delta\rightarrow 0$ then gives the desired estimates for $N_{q}$ ([13], pages 102–103 and [36], section 3.3).
So $q$ is now fixed, $q_{0}\leq q$, and $N_{q+1}$, hence $P_{q}$, satisfy Sobolev estimates. We want to prove Sobolev estimates for $N_{\delta,q}$, with constants uniform in $\delta$ ($\delta$ small). Because $P_{q}$ is regular in Sobolev norms, Lemma 7 in [35] (see also [37], Theorem 1.1) says that for all $s>0$, there is a constant $C_{s}$ independent of $\delta$ such that we have the estimate
@@
For $\eta\in(0,1)$, set $X_{\eta}=e^{\rho_{\eta}g_{\eta}}L_{n}^{\rho_{\eta}}$, where $g_{\eta}$ is a smooth function that agrees on the boundary with $\alpha_{\eta}(\overline{L_{n}^{\rho_{\eta}}})$. Because $\alpha_{\eta}(T^{\rho_{\eta}})=0$, $\alpha_{\eta}\big{(}\overline{L_{n}^{\rho_{\eta}}}\big{)}$ is real, and we may take $g_{\eta}$ to be real valued. This modification ia analogous to replacing $h_{\eta}$ by $\widetilde{h_{\eta}}$ in section 3. Then $(X_{\eta}-\overline{X_{\eta}})$ is tangential at the boundary. The reason for the modification is that the terms where $j=n$ in the commutator (39) below will then vanish on the boundary. This fact is needed in the argument. Being able to make this choice is a manifestation of the principle that ‘commutator conditions in the normal (and any strictly pseudoconvex) direction come for free’ ([36], Section 5.7, [9], proof of the lemma). Note that because $g_{\eta}$ is only prescribed on the boundary, we may assume that $|\rho_{\eta}g_{\eta}|\leq 1/2k$. As usual, we need tangential differential operators to preserve the domain of $\overline{\partial}^{*}$, and so we let them act in special boundary charts ([36], section 2.2). The error this introduces is one order lower than the operator, and as a result, our estimates are not affected.
@@
The $\|u\|_{k}^{2}$ term is needed because of the compactly supported term from the partition of unity used for the special boundary charts; this term is not covered by Lemma 3. The estimate follows from interior elliptic regularity of $\Box_{\delta,q}$, with constants that are uniform in $\delta$. This uniformity holds because for $u\in dom(\Box_{\delta,q})$, $\Box_{\delta,q}u=-(1/4+\delta)\Delta u$, where $\Delta$ acts coefficient wise (see [36], formula (3.22)). We have used that when there is a barred derivative, or a complex tangential derivative, we can always commute it so that it acts first. Then there is no issue with the requirement that a form must be in the domain of $\overline{\partial}^{*}$ in order for Lemma 1 to apply. The error term this makes is of order at most $(k-1)$. Modulo terms controlled by ¯$\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}$, we have $\|(e^{-h_{\eta}}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}\lesssim\|e^{-kh_{\eta}}(L_{n}^{\rho}-\overline{L_{n}^{\rho}})^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}\leq e\|(X_{\eta}-\overline{X_{\eta}})^{k}\overline{\partial}^{*}N_{\delta,q}u\|^{2}$ (since $2k|\rho g_{\eta}|\leq 1$). Upon also estimating the second to last term in (30) via a s.c.–l.c argument and absorbing the term $s.c\|\overline{\partial}^{*}N_{\delta,q}u\|_{k,2h_{\eta}}^{2}$ into the left hand side of (30), we thus arrive at
@@
Here, $\vartheta$ is the formal adjoint of $\overline{\partial}$, $(\partial/\partial\nu)$ denotes the normal derivative acting coefficient wise on forms, and $\omega_{n}$ is the form dual to the unit normal $L_{n}$. The passage from $\|(e^{-h_{\eta}}(L_{n}^{\rho}-\overline{L_{n}^{\rho}}))^{k}\overline{\partial}N_{\delta,q}u\|^{2}$ to $\|(X_{\eta}-\overline{X_{\eta}})^{k}\overline{\partial}N_{\delta,q}u\|^{2}$ is as in (30) – (31).
@@
More precisely, it is shown in [35] that for $\eta$ given, there is $\delta_{0}(\eta)$ such that all the error terms in (31)–(33) are acceptable for (29) (for $l=k$), or can be absorbed into the left hand side of (29) when $\delta\leq\delta_{0}(\eta)$. We will use ‘acceptable’ somewhat loosely for terms in an estimate that are either bounded appropriately by the right hand side, or can be absorbed into the left hand side. Note that upon adding (34) through (36), using Cauchy–Schwarz on each inner product, followed by a s.c.–l.c. argument, the squares of the norms of the left hand sides in (34)–(36) can be absorbed into the first term on the right hand side of the sum of (31)–(33). Therefore, in order to estimate the sum of the left hand sides of (31)–(33), we only have to estimate
@@
For the commutators with $\overline{\partial}$ and $\overline{\partial}^{*}$ we use Lemmas 3 and 4, with $\rho=\rho_{\eta}$, $h=-\rho_{\eta}g_{\eta}$ and $v=N_{\delta,q}u$. It suffices to estimate the terms in (38) with $N_{\delta,q}u$ replaced by $\chi N_{\delta,q}u$, where $\chi\in C^{\infty}(\mathbb{C}^{n})$ is supported in a boundary chart. Via a suitable partition of unity (independent of $\eta$), these estimates will then be summed to obtain global estimates. Lemma 3 gives
@@
(39) [¯$\|[\overline{\partial},(X_{\eta}-\overline{X_{\eta}})^{k}]\chi N_{\delta,q}u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j\notin J,j<n}\int_{\Omega}\big{|}\overline{L_{j}}(\rho_{\eta}g_{\eta})-\alpha_{\eta}(\overline{L_{j}})\big{|}^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\;\;\;\;\;\;\;\;\;\;\;\;\\ +\sideset{}{{}^{\prime}}{\sum}_{n\notin J}\int_{\Omega}\big{|}\overline{L_{n}^{\rho_{\eta}}}(\rho_{\eta}g_{\eta})-\alpha_{\eta}(\overline{L_{n}^{\rho_{\eta}}})\big{|}^{2}|(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{J}|^{2}\;\;\;\;\;\;\;\\ \;\;\;\;+C_{\eta}\,(\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}+\|u\|_{k-2}^{2})\;.$
@@
We have used (28) to estimate the terms of order at most $(k-1)$. The sum is only over $j\notin J$ because for $j\in J$, the term $\overline{\omega_{j}}\wedge\overline{\omega_{J}}$ vanishes. In the error terms, we omit the factor $\chi$ and use that the commutators of $\chi$ with $\overline{\partial}$ and $\overline{\partial}^{*}$ are of order zero. Because there are only finitely many $\chi$s, we may keep $C_{\eta}$ independent of $\chi$. The term $\|u\|_{k-2}^{2}$ again results from the compactly supported term that arises from letting $(X_{\eta}-\overline{X_{\eta}})$ act in special boundary charts; as above, the estimate follows from interior elliptic regularity of $\Box_{\delta,q}$, with constants that are uniform in $\delta$. The order $(k-1)$ error terms are acceptable by the induction hypothesis, and using (28) for $\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}$ shows that this term can be split into an absorbable term and a lower order term. From Lemma 4, we similarly obtain (but not taking the absolute values all the way inside the sums)
@@
(40) [¯$\|[\overline{\partial}^{*},(X_{\eta}-\overline{X_{\eta}})^{k}]\chi N_{\delta,q}u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{S}\int_{\Omega}\big{|}\sum_{j}\big{(}L_{j}(\rho_{\eta}g_{\eta})-\alpha_{\eta}(L_{j})\big{)}(X_{\eta}-\overline{X_{\eta}})^{k}(\chi N_{\delta,q}u)_{jS}\big{|}^{2}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;+C_{\eta}\,\left(\|\overline{\partial}N_{\delta,q}u\|_{k-1}^{2}+\|\overline{\partial}^{*}N_{\delta,q}u\|_{k-1}^{2}+\|N_{\delta,q}u\|_{k-1}\|N_{\delta,q}u\|_{k}+\|u\|_{k-2}^{2}\right)\;.$